OpenStudy (destro):

What property is shown by the equation? Column A Column B 1. (–22 + 17) + 13 = –22 + (17 + 13) 2. –(27 + (–3)) = –27 + 3 3. 35 + (–15 + 4) = (35 + (–15)) + 4 4. –12 + 6 = 6 + (–12) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (midhun.madhu1987):

(a+ b) + c = a + (b + c) - associative property a +b = b + a -- commutative property